OpenStudy (anonymous):

Find the center, foci , asymptotes of the hyperbola y^2 - x^2 - 4y + 10x -70 = 0

OpenStudy (amistre64):

+4+25 = +29+70 = 99 hmmm (y-2)(x+5) center = (-5,2)

OpenStudy (amistre64):

(y-2)^2 (x+5)^2 ------ - ------- = 1 99 99 its a square under there lol.... a and b = sqrt(99) vertex is along the y so (-5, 2+sqrt(99)) and (-5, 2-sqrt(99))

OpenStudy (amistre64):

asymps are y = +-(x+5) +2 right?

OpenStudy (amistre64):

maybe not lol

OpenStudy (amistre64):

its a 45 from x = -5,2 y = +- (x+5)+2 y = x+7 and y = -x-3 maybe

OpenStudy (amistre64):

foci are... what; sqrt(198) from center right? (-5,2-sqrt(189))(-5,2+sqrt(189)) if i did it right lol